Circular No. 7405

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-3 E. B. Waltman, Naval Research Laboratory; and F. D. Ghigo, National

Radio Astronomy Observatory, report: ”The Green Bank Interferometer has observed the strongest flare in Cyg X-3 since 1991. Flux densities reached 17 Jy at 8 GHz and 11 Jy at 2 GHz on Apr. 19.6 UT, following an extended period of very low hard x-rays (IAUC 7365). The data may be

accessed at http://www.gb.nrao.edu.”

SUPERNOVA 1999gv IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, report: “Comparison

between Hubble Space Telescope (HST) WFPC2 images (all with limiting mag 25 or fainter) of the rich galaxy cluster Abell 2218, obtained with the F606W filter on 2000 Jan. 11–13 and 1999 Mar. 11 and with the F702W filter (considered to have a ‘broad B’ bandpass) on 1994 Sept. 3, reveals an unresolved object that is present in images obtained in March 1999 but absent from the more recent and the older images. The transient object (mag 22.8 in the HST F606W filter system, which is considered to have a ‘broad V ’ bandpass) is located 2′′.8 west and 0′′.8 south of the nucleus of one of the brighter cluster galaxies, whose center is located at α = 16^{h}36^{m}00^{s}.74, δ = +66^{o}13′19′′.9 (equinox 2000.0). SN 1999gv is present in twelve consecutive images taken between 1999 Mar. 11.084 and 11.256 UT, and we estimate an upper limit of 0′′.012 for the proper motion of the object during the 4-hr period of observations. In view of the lack of any apparent motion or elongation of the PSF and the close coincidence with a cluster galaxy, we assess that the object is most likely a supernova. The host galaxy is catalogued as number 149 in the list by Le Borgne et al. (1992, A.Ap. Suppl. 95, 87), where its redshift is given as z = 0.1753 and its magnitude as B = 19.89. We are trying to constrain the supernova light curve using groundbased observations of this cluster that we obtained during 1998–1999 with the Keck II 10-m, WIYN 3.5-m, and Wise 1-m telescopes. Observers with relevant data, especially deep images of Abell 2218 taken between Feb. and Apr. 1999, are invited to contact the first author (avishay@wise.tau.ac.il). An image of SN 1999gv can be seen

at http://wise-obs.tau.ac.il/~avishay/a2218.jpg.”

V1333 AQUILAE

Corrigendum. On IAUC 7161, line 12, for 9 hr read 4.3 hr

2000 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green